OpenStudy (anonymous):

The underground water pipe between two streets is damaged. Half the length of the pipe needs to be changed. The layout of the pipes on a coordinate grid shows one end point of the pipe at (-6, 1) and its midpoint at (-2, 3). Troy performed the steps shown to find the coordinates of the other endpoint of the pipe.

OpenStudy (anonymous):

hi yazzy, You seem curious @lgbasallote will help you better than me

OpenStudy (anonymous):

Step 1. x1 = -6, y1 = 1, x = -2, y = 3 Step 2. Step 3. Step 4. 4 = -6 + x2, 6 = 1 + y2 Step 5. 10 = x 2 , 5 = y2 In which step did Troy first make an error? Step 3, because he substituted the incorrect values in the formula. Step 4, because he found the product of -2 and 2 to be 4 instead of -4. Step 2, because he used the incorrect formula to find the midpoint. Step 5, because he found the difference of 6 and 1 to be 5 instead of -5.

OpenStudy (anonymous):

Step 4, because he found the product of -2 and 2 to be 4 instead of -4.

OpenStudy (anonymous):

I m bad @ EXPLAINING

OpenStudy (anonymous):

Omg, thank you so much, can I ask another one ? :)))

OpenStudy (anonymous):

a medal first > other question

OpenStudy (anonymous):

Done. :) An architect drew a figure on a coordinate grid as shown to represent a scaled model of the front view of a garage. What is the length of the segment PQ of the roof rounded to the nearest hundredth? 11.66 units 12.00 units 12.65 units 6.00 units

OpenStudy (anonymous):

as per the rules and law it must be 12.00 units If a fig.. woul be there

OpenStudy (anonymous):

then its 6.00

OpenStudy (anonymous):

How can you check if a segment is an angle bisector of a triangle? Check if measures of the two angles the segment divides the vertex angle into are the same. Check if the areas of the two triangles the segment forms are equal. Check if the segment meets a side of the triangle at a right angle and divides the vertex angle into two congruent angles. Check if the segment meets a side of the triangle at a right angle and divides the side into two congruent parts.

OpenStudy (anonymous):

obviously, Check if the segment meets a side of the triangle at a right angle and divides the side into two congruent parts.

OpenStudy (anonymous):

Thats what I said, and my sister told me it was wrong, omg, you are a life saver!!

OpenStudy (anonymous):

Okay last last last question i promise okay? :D

OpenStudy (anonymous):

thanks I think she asked it to clarify

OpenStudy (anonymous):

all the best

OpenStudy (anonymous):

Haley constructed the obtuse isosceles triangle RST as shown in the figure below. What is the error in Haley’s construction? She plotted point T to the right of the perpendicular line. She did not ensure that side RS is congruent to side RT. She plotted two vertices on the arc. She did not plot point R on the arc.

OpenStudy (anonymous):

fig..

OpenStudy (anonymous):

Look at triangle KLM. Segment ME is 21 inches. Which of these could be the dimensions of the segments in the triangle? ME = 21 inches, MO = 14 inches FK = 20 inches, KO = 10 inches GO = 10 inches, LO = 10 inches MO = 9 inches, EO = 3 inches

OpenStudy (anonymous):

firstly look to it if it's definition " An obtuse angle is more than 90° but less than 180° In other words, it is between a right angle and a straight angle"